OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x2 + 4x - 3 = 0

OpenStudy (unklerhaukus):

\[x^2+4x-3\qquad=0\\x^2+4x\qquad\qquad=3\\x^2+2x+2x\qquad=3\\x^2+2x+2x+2^2\quad=3+2^2\]

OpenStudy (unklerhaukus):

...